OpenStudy (p0sitr0n):

Anyone has interesting Calculus / Abstract Algebra questions or problems ? xD

OpenStudy (anonymous):

Ive got some geometry

OpenStudy (anonymous):

It depends on your definition of "interesting", but give this integral a try: \[\int_0^\infty \frac{\cos2t-\cos t}{t}\,dt\]

OpenStudy (p0sitr0n):

yeah some sort of this, I think I saw it recently on Math exchange :P